# Why We Build Local Large Language Models: An Observational Analysis from 35 Japanese and Multilingual LLMs

### v1

@@ A.2 Evaluation Tasks and Benchmarks
Name Description Lang. Eval. metric1111footnotemark: 11,1212footnotemark: 12 Exp. setup Reference JcommonsenseQA (JCom.) Multiple-choice questions with 5 options based on a knowledge base JA Acc. 4-shot Kurihara et al. (2022) JEMHopQA Free-form question answering to evaluate knowledge and reasoning ability JA Char F1 4-shot Ishii et al. (2023) NIILC Free-form question answering where answers can be obtained from an encyclopedia JA Char F1 4-shot Sekine (2003) JSQuAD Free-form question answering on Wikipedia articles JA Char F1 4-shot Kurihara et al. (2022) XL-Sum Generating summaries from BBC articles JA ROUGE-2 1-shot Hasan et al. (2021) MGSM Japanese translation of the primary school math word problem dataset (GSM8K) JA Acc. (EM) 4-shot Shi et al. (2023) WMT20(en-ja) English-Japanese translation of news articles JA BLEU 4-shot Barrault et al. (2020) WMT20(ja-en) Japanese-to-English translation of news articles JA BLEU 4-shot Barrault et al. (2020) JMMLU Japanese translation of the multiple-choice benchmark MMLU (53 subjects) JA Acc. 5-shot Yin et al. (2024) JHumanEval Japanese translation of HumanEval JA pass@1 0-shot 10 trials Sato et al. (2024) OpenBookQA Multiple-choice questions based on scientific knowledge and common sense EN Acc. 4-shot Mihaylov et al. (2018) TriviaQA Free-form question answering based on trivia knowledge EN Acc. (EM) 4-shot Joshi et al. (2017) HellaSwag Multiple-choice questions to predict the next event EN Acc. 4-shot Zellers et al. (2019) SQuAD2 Free-form question answering based on a supporting document EN Acc. (EM) 4-shot Rajpurkar et al. (2018) XWINO Binary-choice questions to identify the antecedent of a pronoun in a sentence EN Acc. 4-shot Tikhonov and Ryabinin (2021) MMLU Multiple-choice questions across 57 subjects EN Acc. 5-shot Hendrycks et al. (2021) GSM8K Primary school math word problem dataset EN Acc. (EM) 4-shot Cobbe et al. (2021) BBH 23 challenging tasks from the BIG-Bench dataset EN Acc. (EM) 3-shot CoT Suzgun et al. (2023) HumanEval Evaluation of code generation ability via unit tests EN pass@1 0-shot 10 trials Chen et al. (2021)
A.3 Estimating the Number of Training Tokens
@@
A.4 Evaluation Framework
Table 3 reports a list of evaluation frameworks used in this study. The table shows the framework name, a brief description, and the reference of the framework. We slightly customized these evaluation frameworks to cover benchmarks that are not officially supported and to implement workarounds for LLMs; for example, some LLMs require special tokens or line breaks in the prompt to generate valid outputs. We will release the customized implementation upon acceptance.
@@ A.4 Evaluation Framework
Name Description Reference LLM-jp eval (1.3.0) Automatic evaluation tool for Japanese LLMs Han et al. (2024) JP Language Model Evaluation Harness (commit #9b42d41) An evaluation framework for Japanese LLMs zenodo.10256836 Language Model Evaluation Harness (0.4.2) An evaluation framework for LLMs zenodo.10256836 Code Generation LM Evaluation Harness (commit #0261c52) An evaluation framework for code generation task Ben Allal et al. (2022) swallow-evaluation An evaluation framework encompassing all the above-mentioned tools 5
A.5 Details of LLM Grouping
@@
Appendix B Analysis of the Evaluation Results
@@
B.1 Performance Difference between the Pre-trained Languages
@@
B.2 Variations in Task Scores
@@
Appendix C Robustness Check of Findings Obtained from Experimental Results
To test the robustness of the findings presented in § 4, we conducted two additional analyses using different methods and settings: the use of maximum likelihood estimation and Promax rotation [^13] We used the factor_analyzer.FactorAnalyzer() and factor_analyzer.Rotator() method from the factor_analyzer package. instead of PCA (in § 4.3); and exclusion of continually pre-trained models to focus on models trained from scratch.
C.1 Maximum Likelihood Estimation and Promax Rotation
@@
C.2 Analysis with only Full-scratch Models

### v2

@@ B.2 Evaluation Tasks and Benchmarks
Name Description Lang. Eval. metric99footnotemark: 9,1010footnotemark: 10 Exp. setup Reference JcommonsenseQA (JCom.) Multiple-choice questions with 5 options based on a knowledge base JA Acc. 4-shot Kurihara et al. (2022) JEMHopQA Free-form question answering to evaluate knowledge and reasoning ability JA Char F1 4-shot Ishii et al. (2023) NIILC Free-form question answering where answers can be obtained from an encyclopedia JA Char F1 4-shot Sekine (2003) JSQuAD Free-form question answering on Wikipedia articles JA Char F1 4-shot Kurihara et al. (2022) XL-Sum Generating summaries from BBC articles JA ROUGE-2 1-shot Hasan et al. (2021) MGSM Japanese translation of the primary school math word problem dataset (GSM8K) JA Acc. (EM) 4-shot Shi et al. (2023) WMT20(en-ja) English-Japanese translation of news articles JA BLEU 4-shot Barrault et al. (2020) WMT20(ja-en) Japanese-to-English translation of news articles JA BLEU 4-shot Barrault et al. (2020) JMMLU Japanese translation of the multiple-choice benchmark MMLU (53 subjects) JA Acc. 5-shot Yin et al. (2024) JHumanEval Japanese translation of HumanEval JA pass@1 0-shot 10 trials Sato et al. (2024) OpenBookQA Multiple-choice questions based on scientific knowledge and common sense EN Acc. 4-shot Mihaylov et al. (2018) TriviaQA Free-form question answering based on trivia knowledge EN Acc. (EM) 4-shot Joshi et al. (2017) HellaSwag Multiple-choice questions to predict the next event EN Acc. 4-shot Zellers et al. (2019) SQuAD2 Free-form question answering based on a supporting document EN Acc. (EM) 4-shot Rajpurkar et al. (2018) XWINO Binary-choice questions to identify the antecedent of a pronoun in a sentence EN Acc. 4-shot Tikhonov & Ryabinin (2021) MMLU Multiple-choice questions across 57 subjects EN Acc. 5-shot Hendrycks et al. (2021) GSM8K Primary school math word problem dataset EN Acc. (EM) 4-shot Cobbe et al. (2021) BBH 23 challenging tasks from the BIG-Bench dataset EN Acc. (EM) 3-shot CoT Suzgun et al. (2023) HumanEval Evaluation of code generation ability via unit tests EN pass@1 0-shot 10 trials Chen et al. (2021)
B.3 Estimating the Number of Training Tokens
@@
B.4 Evaluation Framework
Table 4 reports a list of evaluation frameworks used in this study. The table shows the framework name, a brief description, and the reference of the framework. We slightly customized these evaluation frameworks to cover benchmarks that are not officially supported and to implement workarounds for LLMs; for example, some LLMs require special tokens or line breaks in the prompt to generate valid outputs.
@@ B.4 Evaluation Framework
Name Description Reference LLM-jp eval (1.3.0) Automatic evaluation tool for Japanese LLMs Han et al. (2024) JP Language Model Evaluation Harness (commit #9b42d41) An evaluation framework for Japanese LLMs zenodo.10256836 Language Model Evaluation Harness (0.4.2) An evaluation framework for LLMs zenodo.10256836 Code Generation LM Evaluation Harness (commit #0261c52) An evaluation framework for code generation task Ben Allal et al. (2022)
B.5 Details of LLM Grouping
@@
Appendix C Analysis of the Evaluation Results
@@
C.1 Performance Difference between the Pre-trained Languages
@@
C.2 Variations in Task Scores
@@
Appendix D Robustness Check of Findings Obtained from Experimental Results
To test the robustness of the findings presented in § 4, we conducted two additional analyses using different methods and settings: the use of maximum likelihood estimation and Promax rotation [^11] We used the factor_analyzer.FactorAnalyzer() and factor_analyzer.Rotator() method from the factor_analyzer package. instead of PCA (in § 4.3); and exclusion of continually pre-trained models to focus on models trained from scratch. Moreover, we performed leave-one-out cross-validation to confirm that our insights derived from observational approach are robust to statistical errors.
D.1 Maximum Likelihood Estimation and Promax Rotation
@@
D.2 Analysis with only Full-scratch Models
@@
D.3 Leave-One-Out Cross-Validation
We assessed the statistical error of factor loadings using leave-one-out cross-validation on the analyzed LLMs (see Figure 17) and confirmed that the standard deviations were small relative to the absolute values.